# Compact Binary Formation in Open Star Clusters II: Difficulty of Gaia NS formation in low-mass star clusters

### v1

@@ 3.2 Effects of CCSN natal kicks
The no-kick model has a high formation efficiency of inert NSBs, in contrast to other models. As described above, a dominant fraction of inert NSBs ($\sim 70$) originate from primordial binaries. We describe their formation pathway here. We do not illustrate it, because we suppose the no-kick model is unrealistic. Let’s prepare a binary with $\sim 8M_{\odot}$ and $\sim 1M_{\odot}$ MSs. It separation is $\sim 3\times 10^{3}R_{\odot}$. The heavier star evolves to a PMS, and fills its Roche lobe. Then, the binary experiences common envelope evolution. The common envelope evolution leaves a binary with $\sim 2M_{\odot}$ He star and $\sim 1M_{\odot}$ MS star. It is separated by $\sim 40R_{\odot}$. The He star causes a CCSN, and leaves behind a NS. Because of CCSN mass loss, the binary is widen, and fits in an inert NSB. In all the models with CCSN natal kicks, such inert NSBs cannot be formed. CCSN natal kicks disrupt the binary at a high probability. Just before CCSNe, the velocities of binary internal motions have $\sim 100$ km s-1, much smaller than CCSN natal kick velocities ($265$ km s-1).
@@
Figure 8: Cumulative eccentricity distribution of inert NSBs in all models except for the no-kick model (Inert NSB w/ kick), and in the no-kick model (Inert NSB w/o kick). The dotted curve shows the thermal distribution.
Figure 8 shows the cumulative eccentricity distribution of inert NSBs in the no-kick model. Their orbital eccentricities appear to deviate from the thermal distribution. In fact, the deviation is significant; the P-value of the K-S test is $2.5\times 10^{-3}$ as seen in Table 2. This is because a dominant fraction of these inert NSBs originate from primordial binaries, not formed through dynamical interactions. On the other hand, the eccentricity distribution of these NSBs is not significantly different from Gaia NS1; the P-value is $5.7\times 10^{-2}$. This shows that it is difficult to distinguish the differences among these eccentricity distributions due to the small statistics.
@@ 4 Discussion
First, we discuss whether inert BHBs and NSBs correspond to Gaia BHs and NSs, respectively. Here, we do not consider the simulation results of the no-kick model. As seen in Table 2, the P-values of the K-S test do not show that inert BHBs have different orbital eccentricities from Gaia BHs. The eccentricity distributions of inert NSBs with $10^{2}\leq P/{\rm day}\leq 10^{4}$ and inert NSBs selected in Table 2 are not inconsistent with Gaia NS1’s eccentricity.
Indeed, only Gaia NS1’s orbital elements have been publicly reported among Gaia NSs, and moreover we do not obtain a sufficient number of inert NSBs from our simulations. Only based on their orbital eccentricities, we cannot robustly conclude if Gaia NS1 is formed in open clusters, or not. Here, we compare the formation efficiencies of inert BHBs and NSBs, assuming that inert NSBs correspond to Gaia NSs. The formation efficiency of inert NSBs is smaller than that of inert BHBs by two or more orders of magnitude except for the no-kick model (see Figure 1). However, the number of discovered Gaia NSs, $20$ (El-Badry et al., 2024), is much larger than the number of Gaia BHs, $2$. This shows that open clusters cannot form enough Gaia NSs to be already discovered.
@@
It is unclear how efficiently Gaia (Gaia Collaboration et al., 2023b) and its spectroscopic follow-up observations (El-Badry et al., 2023b, c, 2024; Chakrabarti et al., 2023) can discover Gaia BHs and NSs. Thus, the numbers of discovered Gaia BHs and NSs do not directly mean their intrinsic numbers. However, for the following reason, we infer that Gaia NSs are harder to be discovered than Gaia BHs, and that the intrinsic number of Gaia NSs should be much larger than that of Gaia BHs. All Gaia and its follow-up observations search for Gaia BHs and NSs, observing BHs’ and NSs’ companions. Because such companions are swayed more largely with BH and NS masses increasing, Gaia, astrometric observations, can more easily discover BHs and NSs with larger masses. Similarly, companions tend to have larger radial velocity variations, and thus spectroscopic follow-up observations can detect more massive BHs and NSs more efficiently. Since NSs are generally less massive than BHs, the intrinsic number ratio of Gaia NSs to Gaia BHs should be larger than the reported number ratio. As seen in section 3, open clusters form Gaia BHs more efficiently than Gaia NSs. This is in contrast to the observational results. Eventually, open clusters cannot form both of Gaia BHs and Gaia NSs.
As described above, Gaia NSs are unlikely to be formed in open clusters. Nevertheless, we do not mean that Gaia BHs are not formed in open clusters. The formation efficiency of Gaia BHs in $\sim 10^{3}M_{\odot}$ open clusters is still high ($\sim 10^{-6}$-$10^{-5}M_{\odot}$), if we do not adopt unrealistically small half-mass density ($\rho\sim 2$ $M_{\odot}~{}{\rm pc}^{-3}$) nor unrealistically small binary fraction ($f_{\rm b,<5M_{\odot}}\sim 0$). It is reasonable that open clusters have $\rho\gtrsim 20$ $M_{\odot}~{}{\rm pc}^{-3}$ initially. Note that, although nearby open clusters have $\rho\sim 10^{-2}$-$10^{2}$ $M_{\odot}~{}{\rm pc}^{-3}$, they should become less dense with time through gas expulsion, stellar evolution mass loss, and two-body relaxation (Portegies Zwart et al., 2010). Initial binary fractions in open clusters should be $f_{\rm b,<5M_{\odot}}\gtrsim 0.2$, similarly to isolated fields (Sana et al., 2012; Moe & Di Stefano, 2017). It would be possible that Gaia BHs are formed in open clusters, while Gaia NSs (El-Badry et al., 2024) and ultramassive WDs (Yamaguchi et al., 2024) are formed on isolated fields. This will become clear after much more Gaia BHs and NSs will be discovered.
@@ 5 Summary
We have performed gravitational $N$-body simulations for $1.3\times 10^{8}M_{\odot}$ open clusters in total with various masses, metallicities, densities, binary fractions, and natal kick velocities. We have found that the formation efficiency is $10^{-6}$-$10^{-5}M_{\odot}^{-1}$ for Gaia BHs in open clusters with realistic parameters, while $\lesssim 10^{-7}M_{\odot}^{-1}$ for Gaia NSs. This is in contrast to observational results in which the number of discovered Gaia BHs ($2$) is smaller than that of Gaia NSs ($\sim 20$). On the other hand, inert BHBs and NSBs obtained from our simulations have orbital eccentricities not inconsistent with Gaia BHs and NSs, respectively. This is partly because of the small statistics for Gaia BHs and NSs, where only one Gaia NS’s orbital elements are publicly available.
@@
Regardless of the presence and absence of CCSN natal kicks, inert NSBs are formed in open clusters less efficiently than, or at most comparable to inert BHBs. Additionally, if there are no CCSN natal kicks, inert NSBs are formed more efficiently on isolated fields than in open clusters. Thus, we have concluded that open clusters cannot form both Gaia BHs and NSs. However, we do not mean that open clusters cannot form Gaia BHs. It would be possible that Gaia BHs are formed in open clusters, while Gaia NSs are formed on isolated fields. We will make it clear when more many Gaia BHs and NSs will be discovered in the future.

### v2

@@ 3.2 Effects of CCSN natal kicks
The no-kick model has a high formation efficiency of inert NSBs, in contrast to other models. As described above, a dominant fraction of inert NSBs ($\sim 70$) originate from primordial binaries. We describe their formation pathway here. We do not illustrate it, because we suppose the no-kick model is unrealistic. Let’s prepare a binary with $\sim 8M_{\odot}$ and $\sim 1M_{\odot}$ MSs. It separation is $\sim 3\times 10^{3}R_{\odot}$. The heavier star evolves to a PMS, and fills its Roche lobe. Then, the binary experiences common envelope evolution. The common envelope evolution leaves a binary with $\sim 2M_{\odot}$ He star and $\sim 1M_{\odot}$ MS star. It is separated by $\sim 40R_{\odot}$. The He star causes a CCSN, and leaves behind a NS. Because of CCSN mass loss, the binary is widened, and fits in an inert NSB. In all the models with CCSN natal kicks, such inert NSBs cannot be formed. CCSN natal kicks disrupt the binary at a high probability. Just before CCSNe, the velocities of binary internal motions have $\sim 100$ km s-1, much smaller than CCSN natal kick velocities ($265$ km s-1).
@@
Figure 8: Cumulative eccentricity distribution of inert NSBs in all models except for the no-kick model (Inert NSB w/ kick), and in the no-kick model (Inert NSB w/o kick). The thick solid curve indicates all the Gaia NSs. The dotted curve shows the thermal distribution.
Figure 8 shows the cumulative eccentricity distribution of inert NSBs in the no-kick model. Their orbital eccentricities appear to deviate from the thermal distribution. In fact, the deviation is significant; the P-value of the K-S test is $2.5\times 10^{-3}$ as seen in Table 2. This is because a dominant fraction of these inert NSBs originate from primordial binaries, not formed through dynamical interactions. On the other hand, the eccentricity distribution of these NSBs is not significantly different from Gaia NS1; the P-value is $5.7\times 10^{-2}$. This shows that it is difficult to distinguish the differences among these eccentricity distributions due to the small statistics if we use only Gaia NS1. However, when we use all the Gaia NSs, the P-value is small, $1.0\times 10^{-5}$. Because of improved statistics, we can conclude that the eccentricity distribution of inert NSBs in the no-kick model is different from that of all Gaia NSs.
@@ 4 Discussion
First, we discuss whether inert BHBs and NSBs correspond to Gaia BHs and NSs, respectively. Here, we do not consider the simulation results of the no-kick model. As seen in Table 2, the P-values of the K-S test do not show that inert BHBs have different orbital eccentricities from Gaia BHs. The eccentricity distributions of inert NSBs with $10^{2}\leq P/{\rm day}\leq 10^{4}$ and inert NSBs selected in Table 2 are not inconsistent with Gaia NS1’s eccentricity. However, when we consider all the Gaia NSs, the eccentricity distribution of inert NSBs is different from that of all Gaia NSs. Thus, it is difficult to identify our inert NSBs as Gaia NSs.
Even if we identify our inert NSBs as Gaia NSs, there is another problem. We compare the formation efficiencies of inert BHBs and NSBs, assuming that inert NSBs correspond to Gaia NSs. The formation efficiency of inert NSBs is smaller than that of inert BHBs by two or more orders of magnitude except for the no-kick model (see Figure 1). However, the number of discovered Gaia NSs, $21$ (El-Badry et al., 2024b, a), is much larger than the number of Gaia BHs, $2$. This shows that open clusters cannot form enough Gaia NSs to be already discovered.
@@
It is unclear how efficiently Gaia (Gaia Collaboration et al., 2023b) and its spectroscopic follow-up observations (El-Badry et al., 2023b, c, 2024b, 2024a; Chakrabarti et al., 2023; Gaia Collaboration et al., 2024) can discover Gaia BHs and NSs. Thus, the numbers of discovered Gaia BHs and NSs do not directly mean their intrinsic numbers. However, for the following reason, we infer that Gaia NSs are harder to be discovered than Gaia BHs, and that the intrinsic number of Gaia NSs should be much larger than that of Gaia BHs. All Gaia and its follow-up observations search for Gaia BHs and NSs, observing BHs’ and NSs’ companions. Because such companions are swayed more largely with BH and NS masses increasing, Gaia, astrometric observations, can more easily discover BHs and NSs with larger masses. Similarly, companions tend to have larger radial velocity variations, and thus spectroscopic follow-up observations can detect more massive BHs and NSs more efficiently. Since NSs are generally less massive than BHs, the intrinsic number ratio of Gaia NSs to Gaia BHs should be larger than the reported number ratio. As seen in section 3, open clusters form Gaia BHs more efficiently than Gaia NSs. This is in contrast to the observational results. Eventually, open clusters cannot form both of Gaia BHs and Gaia NSs.
As described above, Gaia NSs are unlikely to be formed in open clusters. Nevertheless, we do not mean that Gaia BHs are not formed in open clusters. The formation efficiency of Gaia BHs in $\sim 10^{3}M_{\odot}$ open clusters is still high ($\sim 10^{-6}$-$10^{-5}M_{\odot}$), if we do not adopt unrealistically small half-mass density ($\rho\sim 2$ $M_{\odot}~{}{\rm pc}^{-3}$) nor unrealistically small binary fraction ($f_{\rm b,<5M_{\odot}}\sim 0$). It is reasonable that open clusters have $\rho\gtrsim 20$ $M_{\odot}~{}{\rm pc}^{-3}$ initially. Note that, although nearby open clusters have $\rho\sim 10^{-2}$-$10^{2}$ $M_{\odot}~{}{\rm pc}^{-3}$, they should become less dense with time through gas expulsion, stellar evolution mass loss, and two-body relaxation (Portegies Zwart et al., 2010). Initial binary fractions in open clusters should be $f_{\rm b,<5M_{\odot}}\gtrsim 0.2$, similarly to isolated fields (Sana et al., 2012; Moe & Di Stefano, 2017). It would be possible that Gaia BHs are formed in open clusters, while Gaia NSs (El-Badry et al., 2024b) and ultramassive WDs (Yamaguchi et al., 2024) are formed on isolated fields. This will become clear after much more Gaia BHs will be discovered.
@@ 5 Summary
We have performed gravitational $N$-body simulations for $1.3\times 10^{8}M_{\odot}$ open clusters in total with various masses, metallicities, densities, binary fractions, and natal kick velocities. We have found that the formation efficiency is $10^{-6}$-$10^{-5}M_{\odot}^{-1}$ for Gaia BHs in open clusters with realistic parameters, while $\lesssim 10^{-7}M_{\odot}^{-1}$ for Gaia NSs. This is in contrast to observational results in which the number of discovered Gaia BHs ($3$) is smaller than that of Gaia NSs ($21$). Moreover, inert NSBs obtained from our simulations have different orbital eccentricities from Gaia NSs. Note that inert BHBs obtained from our simulations have orbital eccentricities not inconsistent with Gaia BHs. This is partly because of the small statistics for Gaia BHs.
@@
Regardless of the presence and absence of CCSN natal kicks, inert NSBs are formed in open clusters less efficiently than, or at most comparable to inert BHBs. Additionally, if there are no CCSN natal kicks, inert NSBs are formed more efficiently on isolated fields than in open clusters. Thus, we have concluded that open clusters cannot form both Gaia BHs and NSs. However, we do not mean that open clusters cannot form Gaia BHs. It would be possible that Gaia BHs are formed in open clusters, while Gaia NSs are formed on isolated fields. We will make it clear when more many Gaia BHs will be discovered in the future.